# Large Language Models Are Active Critics in NLG Evaluation

### v1

@@ Abstract
The conventional paradigm of using large language models (LLMs) for evaluating natural language generation (NLG) systems typically relies on two key inputs: (1) a clear definition of the NLG task to be evaluated and (2) a list of pre-defined evaluation criteria. This process treats LLMs as “passive critics,” strictly following human-defined criteria for evaluation. However, as new NLG tasks emerge, the criteria for assessing text quality can vary greatly. Consequently, these rigid evaluation methods struggle to adapt to diverse NLG tasks without extensive prompt engineering customized for each specific task. To address this limitation, we introduce Active-Critic, a novel LLM-based NLG evaluation protocol that enables LLMs to function as “active critics.” Specifically, our protocol comprises two key stages. In the first stage, the LLM is instructed to infer the target NLG task and establish relevant evaluation criteria from the data. Building on this self-inferred information, the second stage dynamically optimizes the prompt to guide the LLM toward more human-aligned scoring decisions, while also generating detailed explanations to justify its evaluations. Experiments across four NLG evaluation tasks show that our approach achieves stronger alignment with human judgments than state-of-the-art evaluation methods. Our comprehensive analysis further highlights the effectiveness and explainability of Active-Critic with only a small amount of labeled data. We will share our code and data on GitHub.
@@ 1 Introduction
Recent advances in language technologies have catalyzed the development of natural language generation (NLG) systems, benefiting a variety of downstream applications such as text summarization (Fabbri et al., 2021), dialogue generation (Mehri & Eskenazi, 2020), and storytelling (Guan et al., 2021). However, despite the rapid growth of NLG systems, reliable techniques for automatic NLG evaluation still lay far behind, primarily due to the inherent challenges posed by the open-ended nature of NLG and the diverse demands of different stakeholders. This gap, in return, undermines the dependability of machine-based generators in real-world deployment.
@@
Traditional NLG evaluation methods typically center on a specific criterion and require human-written references for comparison (Li et al., 2024). Commonly considered criteria include similarity-based reference alignment, text fluency, human likeness, and information adequacy. Some NLG tasks, such as text summarization, incorporate additional criteria like coherence and consistency (Lin & Chen, 2023). Notably, the measurement of alignment between machine-generated candidates and human-written references has been widely designed into various metrics, ranging from n-gram matching like BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) to embedding-based matching such as BERTScore (Zhang et al., 2019) and BARTScore (Yuan et al., 2021). To save the efforts of picking multiple appropriate single-aspect evaluators for all criteria regarding a specific NLG task, recent advent of large language models (LLMs) has introduced a new evaluation paradigm that is capable of scoring machine-generated candidates across multiple criteria simultaneously for diverse NLG tasks, either by fine-tuning the LLMs (Zhong et al., 2022; Jiang et al., 2023; Xu et al., 2023; Ke et al., 2023) or by prompting the model for assessment (Chiang & Lee, 2023; Gong & Mao, 2023; Lin et al., 2023). To save the high cost of human annotation on references and avoid potential biases caused by limited references, some latest works further emphasize designing reference-free evaluation methods (Fu et al., 2024; Liu et al., 2023a; Li et al., 2023; Jia et al., 2023).
Despite remarkable contributions made by prior work, existing NLG evaluation approaches commonly require two key inputs: (1) a clear definition of the target NLG task to be evaluated, and (2) a set of predefined evaluation criteria (e.g., coherence, relevance) to guide the machine evaluator’s judgments. Shaped by the developers’ understanding and expectations of the evaluation process, these machine evaluators often function as “passive critics”, constrained by the fixed criteria established by the developers. This limitation may hinder the machine from discovering valuable insights within the data that could enhance its assessment capabilities. Such constraints become even more pronounced when new NLG tasks are introduced, as the criteria for evaluating text quality can vary widely. As a result, these rigid evaluation methods struggle to adapt to diverse NLG tasks without extensive prompt engineering tailored to each specific task.
In this study, we concentrate on the potential of LLMs to evolve into “active critics”, motivated by the impressive performance of recent LLMs in in-context few-shot learning via prompting. We introduce Active-Critic—a novel LLM-based NLG evaluation protocol that assesses NLG systems based solely on the model’s active engagement with data (i.e., machine-generated responses and their human-annotated scores). By automatically inferring what and how to evaluate from a few labeled data, this protocol offers flexibility in assessing diverse NLG tasks, particularly when adapting to new ones. Furthermore, it shows promise in capturing the diverse requirements of different stakeholders through their annotated data, facilitating more personalized evaluations.
To build Active-Critic, we develop a multi-stage evaluation pipeline that directs an LLM to (1) infer the target NLG task and identify relevant evaluation criteria from the data, and (2) optimize prompts to make its scoring better align with human judgments. To enhance the trustworthiness of Active-Critic, we also prompt the model to generate explanations alongside its scoring. Experiments on four NLG tasks using an open- and a closed-source LLM demonstrate that the Active-Critic outperforms various state-of-the-art baselines. Additionally, we show that Active-Critic is helpful in identifying more fine-grained evaluation aspects compared to developer-defined criteria, and its scoring explanations at both the aspect and overall levels can provide valuable insights. In summary, our contributions include:
A novel NLG evaluation protocol: Active-Critic functions as an active critic by automatically inferring the target task, establishing necessary evaluation criteria, refining its scoring, and generating fine-grained explanations to support its decisions.
Extensive experiments across multiple NLG tasks: We conduct extensive experiments to demonstrate the effectiveness and trustworthiness of Active-Critic in diverse NLG tasks.
Comprehensive analysis and findings: We show that combining fine-grained, criteria-specific scoring with explanation generation encourages the LLM to engage more deeply with the test cases, resulting in improved overall quality assessments.
@@ 2 Related Work
NLG Evaluation.
The current landscape of NLG evaluation methods can be viewed from two key perspectives. First, in terms of methodological design, these methods can be categorized into three groups, early human-centric evaluation (Mellish & Dale, 1998), followed by untrained machine evaluation (Papineni et al., 2002; Lin, 2004; Lavie & Denkowski, 2009), and more recently, machine-learned evaluation (Sennrich et al., 2015; Zhang et al., 2019; Yuan et al., 2021; Kim et al., 2023). Second, with respect to functionality, existing studies largely concentrate on single-criteria metric design, targeting either general NLG tasks like reference alignment (Liu et al., 2023b) or a specific NLG task like coherence for text summarization (Wang et al., 2023b). To enhance evaluation efficiency, some latest works have advocated for unified evaluation frameworks built on top of a pre-trained language model, aiming to transcend task-specific boundaries and assess multiple criteria simultaneously (Chiang & Lee, 2023; Liu et al., 2024; Gong & Mao, 2023). Since our work falls into this group, we discuss the details of these works below.
Overall, the study of unified evaluation frameworks encompasses two major strands. The first strand emphasizes the generality of evaluation methods, specifically targeting the estimation of instance quality scores based on defined tasks and criteria, w/wo human-written references (Xiao et al., 2023; Gao et al., 2024). Within this strand, researchers concentrate on two major strategies: (1) developing criteria-centered prompts that guide LLMs for multi-faceted, train-free evaluations (Fu et al., 2024; Liu et al., 2023a; Lin & Chen, 2023), and (2) curating a large-scale multi-scenario benchmark to fine-tune an LLM as a generalized evaluator (Zhong et al., 2022; Li et al., 2023; Wang et al., 2023a; Ke et al., 2023). Comparatively, the first strategy offers a cost-effective approach to evaluation, while the second enhances scoring consistency and reproducibility. Moving beyond generality, the second strand of research focuses on evaluation interpretability, particularly in analyzing prediction errors by prompts (Xu et al., 2023; Jiang et al., 2023). In this study, we address both aspects of evaluation by introducing a new paradigm that guides the model to self-infer the target task and relevant evaluation criteria, ultimately making a final decision with free-text explanations drawn from several rated data instances, enabling the machine to be more active and flexible in evaluation.
@@
Dynamic prompt optimization focuses on iteratively adjusting prompts to improve the performance of static LLMs on specific tasks. Existing methods can be categorized by inference depth into two groups: single-layer and multi-layer prompt optimization. Single-layer methods, such as APE (Zhou et al., 2023), APO (Pryzant et al., 2023), OPRO (Yang et al., 2023), and IPC (Levi et al., 2024), focus on optimizing prompts within a single stage, which limits their adaptability to complex tasks. In contrast, multi-layer methods like DSPy (Khattab et al., 2023) and MIPRO (Opsahl-Ong et al., 2024) optimize prompts across multiple stages, enabling more comprehensive reasoning but relying on scalar-based comparisons between data points, which fall short for tasks requiring correlations across data vectors. Our approach introduces a correlation-based comparison instead of a scalar-based comparison optimizing multi-stage NLG evaluation tasks. Specifically, we redesigned DSPy’s prompt optimization algorithm using a two-stage process of Task Inference and Self-Optimizing Scoring, incorporating correlation-based validation metrics.
3 Preliminary
Problem Definition.
@@ 3 Preliminary
Prior Work.
Existing prompt-based methods using LLMs for evaluation often rely on manually designed rule-based criteria $c_{\text{rule}}$ (e.g., coverage, conciseness) tailored to a particular NLG task (e.g., text summarization), and incorporate these criteria into the prompt as $\text{LLM}(x_{i},y_{i},c_{\text{rule}})\to\hat{r}_{i}$ without explanations. These methods face two major limitations. First, their rule-based criteria may not generalize well across different NLG evaluation tasks. Second, they require human effort to design criteria that align with the considerations human critics use for quality ratings. Consequently, verifying these rule-based criteria against the dataset ${\mathcal{D}}_{\text{train}}$ passively is often considered an afterthought.
@@ 4 Active-Critic
Our Active-Critic is a multi-stage evaluation framework designed for task-adaptive, explainable NLG evaluation. Unlike prior methods that use LLMs as a passive critic, our method optimizes the LLM as an active critic by searching for an optimal evaluation protocol $\Phi^{*}$ from the data itself that maximizes the correlations between the estimated ratings and human ratings in ${\mathcal{D}}_{\text{train}}$.
Specifically, we design two main stages: (1) task inference4.1) that obtains the text description $I$ of the NLG task and its key evaluation criteria; and (2) self-optimizing scoring4.2) that automatically derives the optimal few-shot examples ${\mathcal{D}}_{\text{demo}}$ from ${\mathcal{D}}_{\text{train}}$ for assessment. Figure 1 displays a workflow of these two stages, and Appendix A shows an example of the evaluation protocol $\Phi$ and its corresponding prompt $z$. Particularly, we implement the above optimization problem using DSPy (Khattab et al., 2023), a declarative programming framework, as it offers convenient declarative modules to define each stage of our evaluation process.
Figure 1: Overview of Active-Critic, including two stages: (1) task inference, where the LLM is instructed to derive the target NLG evaluation task description and relevant criteria from data samples, and (2) self-optimizing scoring, allowing the LLM to generate multi-criteria and overall quality scores along with accompanying explanations.
@@ 4.1 Task Inference
The task inference stage, depicted on the left side of Figure 1, focuses on identifying two key components for NLG evaluation: (1) task description and (2) criteria definition. This stage aims to allow the LLM to autonomously analyze the input dataset, infer the characteristics of the NLG task, and establish relevant evaluation criteria without human intervention.
@@
This module focuses on prompting the LLM to formulate an accurate task description $T$ by reviewing examples in $\mathcal{D}_{\text{train}}$ and identifying key information that characterizes the target NLG task (e.g., summarization, storytelling) for evaluation. Considering that LLM’s context length limit may not fit in all examples in ${\mathcal{D}}_{\text{train}}$, we split these examples into $N$ mini-batches, and generate one task description $T_{n}$ from each mini-batch ${\mathcal{D}}_{\text{train},n}$. That is, $\text{LLM}(f_{t}({\mathcal{D}}_{\text{train},n}))\to T_{n},~{}\forall n\in[1,N]$, where $f_{t}$ is a prompt template shown in Table 9 in the appendix. The final task description $T$ is generated by the LLM through the ensemble of all task descriptions $\{T_{n}\}_{n=1}^{N}$ over all mini-batches.
@@
After establishing the task description for each mini-batch, the LLM is instructed to define task-specific evaluation criteria for assessing the quality of machine-generated texts. Unlike traditional evaluation frameworks that rely on predefined criteria (e.g., coherence, fluency), we instruct the LLM to automatically identify the most relevant evaluation dimensions for the target NLG task. Similar to the task description, the final criteria set $C=\{c_{1},c_{2},...,c_{m}\}$ [^2] We instruct the LLM to output a criteria set in the JSON format, as shown in Table 10 in the appendix. is composed of all relevant dimensions inferred from all the reviewed mini-batches.
@@
4.2 Self-Optimizing Scoring
Our second stage, as shown on the right side of Figure 1, focuses on automatically optimizing evaluation prompts to ensure the model delivers detailed, explainable scoring results that closely align with human judgments. Inspired by prior research that harnesses the potential of LLMs by breaking down complex tasks into simpler ones (Wei et al., 2022; Khot et al., 2023), we hypothesize that starting with fine-grained, criteria-specific scoring can help the model derive an accurate overall quality score. With this intuition in mind, we structure the scoring stage into two parts: (1) Multi-criteria Scoring with Explanation (McS-E), followed by (2) Overall Scoring with Explanation (OS-E). The prompts $z$ used in each component are treated as optimization parameters, and these prompts are template-based, generated from the evaluation protocol $\Phi$.

### v2

@@ Abstract
The conventional paradigm of using large language models (LLMs) for natural language generation (NLG) evaluation relies on pre-defined task definitions and evaluation criteria, positioning LLMs as “passive critics” that strictly follow developer-provided guidelines. However, human evaluators often apply implicit criteria, and their expectations in practice can vary widely based on specific end-user needs. Consequently, these rigid evaluation methods struggle to adapt to diverse scenarios without extensive prompt customization. To address this, we introduce Active-Critic, a novel LLM-based evaluator that transforms LLMs into “active critics” capable of adapting to diverse NLG tasks using limited example data. Active-Critic consists of two stages: (1) self-inferring the target NLG task and relevant evaluation criteria, and (2) dynamically optimizing prompts to produce human-aligned scores along with detailed justifications. Our experiments show that Active-Critic can generate nuanced, context-aware evaluation criteria, enabling it to achieve superior alignment with human judgments across multiple tasks.
Large Language Models Are Active Critics in NLG Evaluation
Shuying Xu Tongji University shuyingxu@tongji.edu.cn                        Junjie Hu University of Wisconsin-Madison junjie.hu@wisc.edu                        Ming Jiangthanks: Corresponding author. Indiana University Indianapolis mj200@iu.edu
@@ 1 Introduction
Recent advances in language technologies have accelerated the development of natural language generation (NLG) systems, benefiting a variety of downstream applications such as text summarization (Fabbri et al., 2021), dialogue generation (Mehri and Eskenazi, 2020), and storytelling (Guan et al., 2021). However, despite the rapid progress in NLG systems, reliable techniques for automatic evaluation of NLG systems still lay far behind, primarily due to the inherent challenges posed by the open-ended nature of NLG and the diverse demands of different stakeholders. This gap, in return, undermines the reliability of machine-generated content in real-world applications.
@@
Traditional NLG evaluation methods typically focus on a specific criterion and require human-written references for comparison (Li et al., 2024b). Commonly considered criteria include reference similarity (Papineni et al., 2002; Lin, 2004; Zhang et al., 2019; Yuan et al., 2021), text fluency (Kann et al., 2018; Mutton et al., 2007), human likeness (Song et al., 2025; Jiang et al., 2019), and information adequacy (Adlakha et al., 2024). Moving beyond single-aspect metrics, recent studies propose to use a universal large language model (LLM) as a judge to score machine-generated texts across multiple criteria in diverse NLG tasks, either by fine-tuning  (Zhong et al., 2022; Jiang et al., 2023; Xu et al., 2023; Ke et al., 2023) or by prompting an LLM for assessment (Chiang and Lee, 2023a; Gong and Mao, 2023; Lin et al., 2023). To address the high cost of human annotation and potential biases introduced by limited references, researchers have further developed reference-free LLM-based evaluations (Fu et al., 2024; Liu et al., 2023a; Li et al., 2023; Jia et al., 2023).
Despite the remarkable advancements of prior work, one major concern remains: the reliance on pre-defined evaluation task descriptions and criteria for assessment, forcing LLM evaluators to adhere strictly to developers’ expectations. In contrast, human evaluators often use nuanced, implicit evaluation criteria that extend beyond these pre-defined criteria in practice (Liu et al., 2024b; Clark et al., 2021; Celikyilmaz et al., 2020). While recent studies (Liu et al., 2024b; Li et al., 2024a; Liu et al., 2024a) have explored prompting LLMs to generate evaluation criteria automatically, these methods still rely on pre-defined task descriptions, requiring substantial manual effort to tailor prompts for each NLG task. Moreover, even within the same task, different stakeholders may prioritize distinct evaluation criteria, making it potentially risky to establish a fixed set of criteria in advance without first accounting for human evaluation nuances.
To overcome the above limitations, we propose a novel evaluation approach, i.e., Active-Critic, that instructs an LLM to actively derive an evaluation protocol purely from human-scored data examples. Our approach includes two stages: (1) adaptively inferring the target NLG task and identifying its underlying evaluation criteria that matter most to end users, and (2) dynamically optimizing prompts to produce human-aligned judgments across diverse NLG scenarios. To enhance trustworthiness, Active-Critic also generates detailed text justifications alongside its scoring.
We have conducted experiments across diverse NLG tasks using various base LLMs. The results show that the Active-Critic consistently achieves a noticeably higher correlation with human judgments, indicating its ability to adapt effectively to different NLG evaluation tasks according to different evaluation criteria. Our approach requires as few as 5 human-scored data to obtain a strong correlation with humans, with performance steadily improving as the dataset grows. Further analysis highlights that the task inference stage contributes more to Active-Critic’s performance than the scoring stage, and Active-Critic can effectively identify nuanced, context-aware criteria beyond pre-defined ones. In summary, our method offers three key benefits:
  • Self-adaptive evaluation. Active-Critic can infer any NLG evaluation task, recover human judgment criteria, and make justified assessments directly from data, eliminating the need for pre-defined task descriptions, fixed evaluation criteria, or manual prompt engineering.
  • Accurate judgment alignment. Our two-stage design guides LLMs to mimic human judgment step by step, yielding interpretable justifications while achieving state-of-the-art alignment with human assessments against strong baselines.
  • Generic for diverse LLMs and NLG tasks. Our method operates independently of specific LLMs and evaluation tasks. Our results on four LLM backbones across four NLG tasks showcase its broad applicability.
@@ 2 Related Work
NLG Evaluation Overview.
Existing methods for NLG evaluation span three major strands, including early human-centric evaluation (Mellish and Dale, 1998), followed by untrained machine evaluation (Papineni et al., 2002; Lin, 2004; Lavie and Denkowski, 2009), and more recently, machine-learned evaluation (Sennrich et al., 2015; Zhang et al., 2019; Yuan et al., 2021; Kim et al., 2023). These studies largely concentrate on single-criteria metric design, targeting either general NLG tasks like reference alignment (Liu et al., 2023b) or a specific NLG task like coherence for text summarization (Wang et al., 2023b). To enhance evaluation efficiency, recent works have advocated for unified evaluation frameworks built upon LLMs, aiming to transcend task-specific boundaries and assess multiple criteria simultaneously (Chiang and Lee, 2023a; Liu et al., 2024a; Gong and Mao, 2023; Liu et al., 2024b; Li et al., 2024a). Our work falls into this group, and will discuss the details below.
LLM-based NLG Evaluation.
Prior studies on unified evaluation frameworks primarily focus on enhancing evaluation generalizability, with an emphasis on estimating instance quality scores across various NLG tasks and multiple criteria simutaneously (Xiao et al., 2023; Gao et al., 2024). Studies in this area typically involve two strategies. One is developing criteria-centered prompts that guide LLMs as a judge for multi-faceted, train-free evaluations (Fu et al., 2024; Liu et al., 2023a; Lin and Chen, 2023; Chiang and Lee, 2023b; Li et al., 2024a; Liu et al., 2024b; Yuan et al., 2023). The other focuses on curating a large-scale multi-scenario benchmark to fine-tune an LLM as a generalized evaluator (Zhong et al., 2022; Li et al., 2023; Wang et al., 2023a; Ke et al., 2023; Kim et al., 2023; Hu et al., 2024).
While prompting-based methods are more cost-effective than tuning-based ones, one major concern with these approaches is the sensitivity of LLMs to manual prompts, possibly causing evaluation biases. To address this issue, several latest works have explored instructing LLMs to generate evaluation criteria  (Liu et al., 2024b; Li et al., 2024a) or scoring rubrics (Liu et al., 2024a) based on pre-defined context like the target NLG task description. In essence, criteria generation in these studies implicitly assumes that each NLG task has a fixed set of evaluation criteria. In contrast, we argue that different end-user needs may lead to varying emphases, even for the same NLG task, resulting in criterion and/or rubric variation. To address this, our approach takes a data-driven perspective, instructing the LLM for NLG evaluations through self-inference of all relevant contexts.
@@
Dynamic prompt optimization iteratively refines prompts to enhance the performance of static LLMs on specific tasks. Existing methods can be divided into two categories based on their inference depth. Single-layer optimization methods, such as APE (Zhou et al., 2023), APO (Pryzant et al., 2023), OPRO (Yang et al., 2023), and IPC (Levi et al., 2024), focus on optimizing prompts within a single stage, limiting their adaptability to complex tasks. In contrast, multi-layer optimization methods, like DSPy (Khattab et al., 2023) and MIPRO (Opsahl-Ong et al., 2024), refine prompts across multiple stages, supporting more comprehensive reasoning but relying on scalar-based comparisons between data points, which are insufficient for tasks requiring correlations across data vectors. We design a correlation-based comparison rather than a scalar-based one to optimize multi-stage NLG evaluation tasks.
3 Notations and Problem Definition
@@ 4 Active-Critic
Figure 1 shows the overall workflow of Active-Critic. With the motivation that an ideal unified evaluation framework should flexibly uncover the nuanced evaluation criteria of end users across diverse generation scenarios and make human-aligned judgments, we design a data-driven evaluation framework structured in two stages. The first stage is task inference4.1), where we instruct an LLM to predict task-related information by actively reviewing a small set of human-rated data examples. Through this analysis of the human-rated data, we expect the model to self-infer the details of the target evaluation task and the implicit criteria used by human annotators. The second stage is scoring alignment4.2), where we aim to align the LLM evaluator with human scoring based on the predicted evaluation criteria. Specifically, we design a dynamic prompt optimization method to automatically select the optimal few-shot examples, ${\mathcal{D}}_{\text{demo}}$, from $\mathcal{D}$ which enables the LLM evaluator to achieve human-aligned scoring through in-context prediction.
Figure 1: Overview of Active-Critic, including two stages: (1) task inference, where the LLM is instructed to derive the target NLG evaluation task description and relevant criteria from data samples, and (2) scoring alignment, allowing the LLM to generate multi-criteria and overall quality scores along with accompanying explanations.
@@ 4.1 Task Inference
The task inference stage, depicted on the left side of Figure 1, focuses on identifying two key components for NLG evaluation: (1) task description and (2) criteria definition. This stage uses the LLM to analyze the dataset $\mathcal{D}_{\text{train}}$, infer the characteristics of the NLG task, and establish relevant evaluation criteria without human intervention.
@@
This module instructs the LLM to formulate an accurate task description $T$ by reviewing examples in $\mathcal{D}_{\text{train}}$ and identifying key information that characterizes the target NLG task (e.g., summarization, storytelling) for evaluation. Considering that LLM’s context length limit may not fit in all examples in ${\mathcal{D}}_{\text{train}}$, we split these examples into $N$ mini-batches, and generate one task description $T_{n}$ from each mini-batch ${\mathcal{D}}_{\text{train},n}$. That is, $\text{LLM}(f_{t}({\mathcal{D}}_{\text{train},n}))\to T_{n},~{}\forall n\in[1,N]$, where $f_{t}$ is a prompt template shown in Table 10 in the appendix. The final task description $T$ is generated by the LLM through the ensemble of all task descriptions $\{T_{n}\}_{n=1}^{N}$ over all mini-batches.
@@
After establishing the task description for each mini-batch, the LLM is instructed to define task-specific evaluation criteria for assessing the quality of machine-generated texts. Unlike traditional evaluation frameworks that rely on predefined criteria (e.g., coherence, fluency), we instruct the LLM to automatically identify the most relevant evaluation dimensions for the target NLG task. Similar to the task description, the final criteria set $\hat{C}=\{\hat{c}_{1},\hat{c}_{2},\dots,\hat{c}_{m}\}$ [^1] We instruct the LLM to output a criteria set in the JSON format, as shown in Table 11 in the appendix. is composed of all relevant dimensions inferred from all the reviewed mini-batches.
@@
4.2 Scoring Alignment
Our second stage, as shown on the right side of Figure 1, focuses on aligning the LLM evaluator with human scoring judgments by automatically optimizing the evaluation prompts. Inspired by prior research that harnesses the potential of LLMs by breaking down complex tasks into simpler ones (Wei et al., 2022; Khot et al., 2023), we hypothesize that starting with fine-grained, criteria-specific scoring can help the model further derive an accurate overall quality score. With this intuition in mind, we structure the scoring stage into two modules: (1) Multi-criteria Scoring with Explanation (McS-E), followed by (2) Overall Scoring with Explanation (OS-E).